Restart Night #8 UT 2018-09-09

OSA: D.G. Huerta (LBTO)

Observers/ISA: B. Rothberg, J. Hill (Tucson)

Software: S. Hooper/P. Kubanek (on-call)

ISp: J. Morris

Mtn Manager: R. Hansen

Details

02:10 Some clouds drifting in from the North, so we've given time to J. Morris to work on MODS.

Open

UT 03:52 - pointing check done, moving to colimation star near field (NGC 7027)

UT 04:01 - seeing looks fantastic, switching to do AGW Transforms first, Stone L1

UT 04:05 - got to field, configuring LUCIs for N375, N30FieldStop, FeII, 2.51 sec exposure

luci1.20180909.0514-0515, luci2.20180909.0514-0515 2.51 sec dark images

UT 04:11 - starting Stone_L1 transform collection both sides

UT 04:21 - it used 2 exposures instead of one (last time I took an image I took 2)

516,517 (L1) are the same

518 I took 1 exposure while transform was doing stuff to clear out the NEXP=2 in memory

UT 04:18 CTRL-C'd because I have lost track of whats going on as to why L1 takes 2x images and LUCI-2 takes 1x

-taking a test image on each side to just clear things up

L1 0523, L2 0522 - junk images to set it up so that DIT=2.51, NDIT=1 NEXP=1

LUCI1/2 AGW Transformation Data Collection

UT 04:20 - starting transform fresh, first images in series are:

L1 0524, L2 0523

UT 04:40 GS167 Failed on SX, restarting on #10 (GS 101) GS 112 is L1 #0533

UT 04:45 - can't seem to get back SX GS0, restarting GCS, guider homed (glitch in AGw stage ??) - retrying from #11 which is GS52 (L1 is #0534)

UT 05:01 - SX GS159 failed to move to position:

IRCL: preset, Stone_L1, MODE=GUIDE, GS=159, TS=0, PA=-135.000 (iif_preset)

===> Error in irc_presetTelescope
err = Resmsg[0] = PresetTelescope result status: Error
err = Resmsg[1] = Stop{FLAO}[235]
err = Resmsg[2] = AGw command status: AGW4: Error moving the motor(s)
err = Resmsg[3] = error: setProbePosition: failed moving probe to position (SFP) (x,y,z): -105.571, 143.749, 15.6374 mm. Patrol field limits ?
err = Resmsg[4] = guidestar list successfully changed but failed moving probe to position.

luci1.20180909.0540 is likely bad - either missed GS or some other error, but the image does not match the field

UT 05:02 - restarting on SX skipping GS159 - starting #19

===> Error in irc_presetTelescope
err = Resmsg[0] = PresetTelescope result status: Error
err = Resmsg[1] = Stop{FLAO}[236]
err = Resmsg[2] = AGw command status: AGW4: Error moving the motor(s)
err = Resmsg[3] = isPosAccessible: probe cannot move to requested position (SFP) (x,y): -0.0362426, -0.267427 mm.
err = Resmsg[4] = guidestar list successfully changed but failed moving probe to position.

UT 05:04 - homing AGW4, restarting GCS

UT 05:06 - restarting #19 (GS 126) - first exposure will be luci1.20180909.0541

UT 05:08 - not sure if found GS0 correctly

UT 05:09 - missed GS0 #541 doesn't look like rest of images

UT 05:10 - resending preset left in acquire

IDL> iif_preset,Stone_L1,mode=acquire,gs=0,pa=-135,/left
IDL> iif_preset,/left
UT 05:12 - didn't land on the correct star - recentering

UT 05:16 - LBCs came up while I wasn't paying attention

Voltage Log for Blue Trackers on successful start:

2018/09/09 04:16:04.998000 S B TRACKERS CCDCTRL highClock voltage line 1 is +3.11V ( +3.0V+-1.0V)
2018/09/09 04:16:05.013000 S B TRACKERS CCDCTRL highClock voltage line 2 is +3.02V ( +3.0V+-1.0V)
2018/09/09 04:16:05.045000 S B TRACKERS CCDCTRL highClock voltage line 3 is +3.02V ( +3.0V+-1.0V)
2018/09/09 04:16:05.060000 S B TRACKERS CCDCTRL highClock voltage line 4 is +3.04V ( +3.0V+-1.0V)
2018/09/09 04:16:05.092000 S B TRACKERS CCDCTRL highClock voltage line 5 is -0.41V ( +0.0V+-1.0V)
2018/09/09 04:16:05.107000 S B TRACKERS CCDCTRL highClock voltage line 6 is +0.13V ( +0.0V+-1.0V)
2018/09/09 04:16:05.138000 S B TRACKERS CCDCTRL highClock voltage line 7 is +2.78V ( +3.0V+-1.0V)
2018/09/09 04:16:05.170000 S B TRACKERS CCDCTRL highClock voltage line 8 is +2.82V ( +3.0V+-1.0V)
2018/09/09 04:16:05.185000 S B TRACKERS CCDCTRL highClock voltage line 9 is +2.87V ( +3.0V+-1.0V)
2018/09/09 04:16:05.217000 S B TRACKERS CCDCTRL highClock voltage line 10 is +2.88V ( +3.0V+-1.0V)
2018/09/09 04:16:05.232000 S B TRACKERS CCDCTRL highClock voltage line 11 is +2.10V ( +2.0V+-1.0V)
2018/09/09 04:16:05.263000 S B TRACKERS CCDCTRL highClock voltage line 12 is +2.74V ( +3.0V+-1.0V)
2018/09/09 04:16:05.295000 S B TRACKERS CCDCTRL lowClock voltage line 1 is -8.97V ( -9.0V+-1.0V)
2018/09/09 04:16:05.310000 S B TRACKERS CCDCTRL lowClock voltage line 2 is -8.94V ( -9.0V+-1.0V)
2018/09/09 04:16:05.342000 S B TRACKERS CCDCTRL lowClock voltage line 3 is -9.05V ( -9.0V+-1.0V)
2018/09/09 04:16:05.357000 S B TRACKERS CCDCTRL lowClock voltage line 4 is -9.13V ( -9.0V+-1.0V)
2018/09/09 04:16:05.388000 S B TRACKERS CCDCTRL lowClock voltage line 5 is -0.42V ( +0.0V+-1.0V)
2018/09/09 04:16:05.404000 S B TRACKERS CCDCTRL lowClock voltage line 6 is +0.04V ( +0.0V+-1.0V)
2018/09/09 04:16:05.435000 S B TRACKERS CCDCTRL lowClock voltage line 7 is -8.99V ( -9.0V+-1.0V)
2018/09/09 04:16:05.467000 S B TRACKERS CCDCTRL lowClock voltage line 8 is -9.03V ( -9.0V+-1.0V)
2018/09/09 04:16:05.482000 S B TRACKERS CCDCTRL lowClock voltage line 9 is -9.05V ( -9.0V+-1.0V)
2018/09/09 04:16:05.513000 S B TRACKERS CCDCTRL lowClock voltage line 10 is -9.02V ( -9.0V+-1.0V)
2018/09/09 04:16:05.545000 S B TRACKERS CCDCTRL lowClock voltage line 11 is -7.84V ( -8.0V+-1.0V)
2018/09/09 04:16:05.560000 S B TRACKERS CCDCTRL lowClock voltage line 12 is -8.90V ( -9.0V+-1.0V)

UT 05:19 - still can't find GS on SX side, DX side is about to finish

UT 05:21 - success! GS0 found, restarting again Stone L on SX GS126 - first LUCI-1 image to use is luci1.20180909.0542

UT 05:33 - failure again

IRCL: preset, Stone_L1, MODE=GUIDE, GS=77, TS=0, PA=-135.000 (iif_preset)

===> Error in irc_presetTelescope
err = Resmsg[0] = PresetTelescope result status: Error
err = Resmsg[1] = Stop{FLAO}[250]
err = Resmsg[2] = AGw command status: AGW4: Error moving the motor(s)
err = Resmsg[3] = error: setProbePosition: failed moving probe to position (SFP) (x,y,z): -115.485, 54.9556, 23.1287 mm. Patrol field limits ?
err = Resmsg[4] = guidestar list successfully changed but failed moving probe to position.

UT 05:35 - homing AGW4: home u -4 m -3

UT 05:36 - Restart #26 GS 162

UT 05:52 - SX failed again on last one, we'll punt it for now

IRCL: preset, Stone_L1, MODE=GUIDE, GS=93, TS=0, PA=-135.000 (iif_preset)

===> Error in irc_presetTelescope
err = Resmsg[0] = PresetTelescope result status: Error
err = Resmsg[1] = Stop{FLAO}[264]
err = Resmsg[2] = AGw command status: AGW4: Error moving the motor(s)
err = Resmsg[3] = error: setProbePosition: failed moving probe to position (SFP) (x,y,z): -131.413, 39.6693, 21.9271 mm. Patrol field limits ?
err = Resmsg[4] = guidestar list successfully changed but failed moving probe to position.

(L1 0524-0539 GOOD, 0542-0547, 0548-0556)

L2 data set is: luci2.20180909.0523-0560

L1 data set is: luci1.20180909.0524-0539,0542-0556

UT 05:53 - homing AGW4 again for good measure

UT 05:55 - resending preset rotated 180 degrees to check on pupils in SX

UT 05:58 - collimating SX with IDl for a bit

Transformation Results DLM

UT 06:00 John changed the DX WFScam_hotspot_x,y position from 376,263 to 371.4,259.0, but the result wasn't as expected so he put it back a few minutes later.

LUCI-BINO NS

UT 06:08 - sending preset for Comet 64P/Swift-Gehrels LUCI-BINO NS imaging tests

UT 06:10 - mask error LUCI-2 - N30 FieldStop mask had problem being put back into storage

UT 06:19 - authorizing LUCI-1 only - NS hijacked LUCI-1

-changed wfs exposures to 15sec

UT 06:33 LUCI-2 MOS error fixed by Dave, problem was cabinet was upside down and not enough power in motor to get mask back into slot.

UT 06:35 - H-band looks good, nice elongated stars, nice big coma, K-band looks wonky, not sure why

UT 06:38 - luci1.20180909.0565 shows double stars - guider must have jumped

UT 06:46 - RTD showing weird problems. images look terrible or offset subtraction, but GEIRS is OK - restarted LUCI Observer GUI, seems better behaved

UT 06:56 - Script completed, looks like NS guiding works well

luci1.20180909.0557-0576 (K,H,J,OH1060) - Comet 64P/Swift-Gehrels

NGC 7027 doing all LUCI filters

UT 06:59 - pointing/collimation check for NGC 7027 script - plan is to run through ALL filters both LUCIs.

UT 07:07 - sending preset for NGC 7027 mega-script

UT 07:38 - switched script so that offsets are +/-60" to avoid crud on LUCI-2 top

UT 08:00 - smudge on the P-gamma filter on LUCI-1

UT 08:24 - Completed 8 filters on both LUCIs for NGC 7027 (K, H, J, Br-gamma, OH 1060, P-gamma, P-beta, OH 1190) - good test of WFS and collimation over 90min - seeing < 1"

L1 - 0577-0616 ; L2 0561-0600

Longslit acquisition test for NGC 315

Longslit acquisition test - Telluric and Target G210 H-band 0.5" slit 1.65microns

UT 08:25 - moving to pointing check

UT 08:29 - sending preset to Telluric target HD 9711

UT 08;31 - SX guider repeatedly goes out of focus each time move to new target

UT 08:39 - filter unknown on LUCI-2 - initialized it with Eng Panel, it was on Br Gamma not HeI

UT 08:46 - moving to NGC 315 G210 H-band 0.5" slit 1.65 microns - elevation 85

UT 08:52 - LUCI is waiting for IIF on left side, stopped and started telescope service

UT 08:53 - pressed GO because inserted PAUSE for first line in script, instead of continuing to line 2 it re-sent the preset

UT 08:54 - DX missed guidestar, resending again

UT 09:00 - position has X Y average remove problem WFS SX (image 174 wfsc) - first time tonight (Blame it on Barry!)

UT 09:03 - SX WFS jumped, but we didn't move - Barry is doing alignment

UT 09:09 - starting spectroscopy

UT 09:10 - error committing grating setup LUCI-1 - manually entering this in the IMGUI, 3x failed to set, used eng gui, settled at 1.65 then went back to 1.00 micron - had to abort exposure

UT 09:15 - set to G200 1.93, moving back to G210 1.65 - this seemed to work, LUCI-1 short one exposure

UT 09:27 - Z9,Z10 turned off on SX

UT 09:35 - Script complete L1 0617-0628 ; L2 0601-0613

UT 09:36 - J. Hill clearing active optics on primary to try and get rid of astigmatism and trefoil. See comment on last nights wiki about the interaction of defocus, trefoil, guiding and poor hotspot centering.

MODS AGw wake-up

UT 09:37 - Switching to MODS to check on AGWs - putting LUCIs in safe config

UT 09:46 - MODS1 & 2 opened hatches homed AGWs

UT 10:21 - MODS AGWs collimated on-axis, works fine, except IDL and GCS don't agree about z10 and z11 on MODS-1. These discrepancies were corrected the following night by switching IDL LDG back to the 20121213 normal reconstructor. Also, MODS-1 still shows a drop in photons compared to MODS-2. (IT 6645)

LBC

UT 10:22 - Reconfiguring for LBCs

UT 10:40 - moving to GD50 CP field for collimation

UT 10:43 - running DOHYBRID

UT 10:54 - collimated, doing Co-Pointing

UT 10:57 - starting GD50 all filters

UT 11:08 - GD50 saturated, changed exptimes on Uspec to 4sec, U to 5sec, B to 2sec, note that V-Bessell is 1 sec on LBC-R but its saturated

UT 11:09 - Restarting with V2script with shorter exptimes airmass 1.22

Start with these files:

1109827, 110918

UT 11:29 - GD50 complete, moving to G191B2b, skipping co-point just going to collimate and start taking data

UT 11:31 - collimating on G191B2b field

UT 11:37 - starting G191B2b script airmass 1.11

UT 11:56 - done with G191B2b - closing up - i think clouds are going to roll through my skyflat, will take LUCI cals before handing over to Dave

Close

UT 12:04 - Taking 10 LBC Biases (10 per LBC), starting LUCI-Bino Cals

UT 12:44 - LBCs powered off

UT 12:46 - Finished LUCI cals - putting LUCI in safe config.

-- %USERSIG{JohnHill - 2018-09-09}%
Topic revision: r7 - 10 Sep 2018, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback